calculate de-Broglie wavelength of an electron travelling at 1% of the speed of light.


a) 2.73 x 10-24


b) 2.42 x 10-10


c) 242.2 x 1010


d) none of these


 


plz solve and explain the solution.....

Grade:11

2 Answers

AJAY LOGESH
13 years ago

 

 

                    λ = h/mc 

         In your problem, velocity = c/100 ( since the velocity is 1 %)

                     Sub c=c/100 in above equation, You will get λ = 100 h/mc

                                     Sub values of h,m and c. Arrange the powers of 10. The answer will be 2.42 x 10-10.

                  So, Your answer will be (b) 2042 x 10-10.  Hope you understand. Do well.

vikas askiitian expert
13 years ago

(lamda) = h/mv              (h is planks constant , h = 6.6*10-34 & m is mass of particle)

 v = 1% of speed of light = C/100                 (C is speed of light & C=3*108)

lamda = h/mv

          =100h/mC                                 (m = 9.1*10-31kg , mass of electron)

 

        = 0.241*10-9 m

       =0.241nm      or 2.41(angstorm)